4. Find the sum and product of the eigenvalues of the matrix

2 3 -2
-2 1 1
1 0 2
1
Expert's answer
2021-09-06T13:09:53-0400

Given matrix is:"\begin{bmatrix}
 2 & 3 & -2 \\
 -2 & 1 & 1\\
1 & 0 & 2
\end{bmatrix}"

Now, solving "\begin{vmatrix}
 2-\lambda & 3 & -2 \\
 -2 & 1-\lambda & 1\\
1 & 0 & 2-\lambda
\end{vmatrix}=0" , we get:

"\Rightarrow(2-\lambda)[(1-\lambda)(2-\lambda)-0]-3[-2(2-\lambda)-1]-2[0-1(1-\lambda)]=0\\
\Rightarrow (1-\lambda)(2-\lambda)^2+12-6\lambda+3+2-2\lambda=0\\
\Rightarrow (1-\lambda)(2-\lambda)^2+17-8\lambda=0\\
\Rightarrow 4+\lambda^2-4\lambda-4\lambda-\lambda^3+4\lambda+17-8\lambda=0\\
\Rightarrow -\lambda^3+\lambda^2-12\lambda+21=0\\
\Rightarrow \lambda^3-\lambda^2+12\lambda-21=0\\"

So, sum of eigenvalues of matrix is: "-(\frac{-1}{1})=1"

and product of eigenvalues of matrix is: "-(\frac{21}{1})=-21"